# Quantum Algorithms for the Pathwise Lasso

### v1

@@ 3.1 Lasso solution
It is known [Tib13, Lemma 14] that there is a Lasso solution such that $|\mathcal{A}|\leq\min\{n,d\}$. If the Lasso solution is unique, then the active set has size at most $\min\{n,d\}$. The sufficient condition $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$ for uniqueness has appeared several times in the literature [OPT00b, Fuc05, Wai09, CP09, Tib13]. Since the active set $\mathcal{A}$ depends on the Lasso solution at $\mathbf{y},\mathbf{X},\lambda$, the condition that $\mathbf{X}_{\mathcal{A}}$ has full rank is somewhat circular. Because of this, more natural conditions are assumed which imply $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$, e.g. the general position property [Ros04, Don06, Dos12, Tib13] and the matrix $\mathbf{X}$ being drawn from a continuous probability distribution [Tib13]. We say that a matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ has columns in general position if no $k$-dimensional subspace in $\mathbb{R}^{n}$, for $k<\min\{n,d\}$, contains more than $k+1$ elements from the set {±$\{\pm\mathbf{X}_{1},\dots,\pm\mathbf{X}_{d}\}$, excluding antipodal pairs. It is known [Ros04, Don06, Dos12, Tib13] that this is enough to guarantee the uniqueness of the Lasso solution.
Fact 18 ([Tib13, Lemma 3]).
If the columns of $\mathbf{X}\in\mathbb{R}^{n\times d}$ are in general position, then the Lasso solution is unique for any $\mathbf{y}\in\mathbb{R}^{n}$ and $\lambda>0$ and is given by Eq. (10).
Another sufficient condition was given by Tibshirani [Tib13], who proved the following.
Fact 19 ([Tib13, Lemma 4]).
If the entries of $\mathbf{X}\in\mathbb{R}^{n\times d}$ are drawn from a continuous probability distribution on $\mathbb{R}^{nd}$, then the Lasso solution is unique for any $\mathbf{y}\in\mathbb{R}^{n}$ and $\lambda>0$ and is given by Eq. (10) with probability $1$.
It is well known that, when $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$, the KKT conditions from Fact 16 imply that the regularisation path $\mathcal{P}\triangleq\{\hat{\bm{\beta}}(\lambda)\in\mathbb{R}^{d}:\lambda>0\}$ comprising all the solutions for $\lambda>0$ is continuous piecewise linear, which was first proven by Efron, Hastie, Johnstone, and Tibshirani [EHJT04], confer Eq. (3). We note that, according to Eq. (8), $\bm{\theta}_{t}=(\mathbf{X}_{\mathcal{A}_{t}}^{\top}\mathbf{X}_{\mathcal{A}_{t}})^{+}\bm{\eta}_{\mathcal{A}_{t}}$ for some active set $\mathcal{A}_{t}$. This fact was later extended to the case when $\operatorname{rank}(\mathbf{X}_{\mathcal{A}})<|\mathcal{A}|$ by Tibshirani [Tib13], who proved that the path of solutions with $\mathbf{b}=\mathbf{0}$ in Eq. (8) is continuous and piecewise linear.
Fact 20 ([Tib13, Appendix A]).
@@
[…]
@@
We call kink a point where the direction $\partial\hat{\bm{\beta}}(\lambda)/\partial\lambda$ changes. The path $\{\hat{\bm{\beta}}(\lambda):\lambda_{t+1}<\lambda\leq\lambda_{t}\}$ between two consecutive kinks $\lambda_{t+1}$ and $\lambda_{t}$ thus defines a linear segment. An important consequence of the above result is that any continuous piecewise linear regularisation path has at most $3^{d}$ linear segments, since two different linear segments have different equicorrelation signs $\bm{\eta}\in\{-1,0,1\}^{d}$. This result was improved slightly to $(3^{d}+1)/2$ iterations by Mairal and Yu [MY12], who also exhibited a worst-case Lasso problem with exactly $(3^{d}+1)/2$ number of linear segments.
@@ 3.2 The LARS algorithm
The Lasso solution path $\{\hat{\bm{\beta}}(\lambda):\lambda>0\}$ can be computed by the Least Angle Regression (LARS) algorithm [^2] By LARS algorithm we mean the version of the algorithm that computes the Lasso path and not the version that performs a kind of forward variable selection. which was proposed and named by Efron et al. [EHJT04], although similar ideas have already appeared in the work of Osborne et al. [OPT00b, OPT00a]. Since its introduction, the LARS algorithm have been improved and generalised [RZ07, MY12, Tib13, ${}^{+}$18]. In compressed sensing literature, this algorithm is better known as the Homotopy method [FR13].
Starting at $\lambda=\infty$, where the Lasso solution is trivially $\mathbf{0}\in\mathbb{R}^{d}$, the LARS algorithm iteratively computes all the kinks in the Lasso solution path by decreasing the parameter $\lambda$ and checking for points where the path’s linear trajectory changes. This is done by making sure that the KKT conditions from Fact 16 are always satisfied. When a kink is reached, a coordinate from the solution $\hat{\bm{\beta}}(\lambda)$ will go from non-zero (being in the active set $\mathcal{A}$) to zero (join the inactive set $\mathcal{I}$), or vice-versa, i.e., from zero (leave $\mathcal{I}$) to non-zero (into $\mathcal{A}$). We shall now describe the LARS algorithm in more details, following [Tib13]. Algorithm 1 summarises the LARS algorithm.
The LARS algorithm starts, without loss of generality, at the first kink $\lambda_{0}=\|\mathbf{X}^{\top}\mathbf{y}\|_{\infty}$, since $\hat{\bm{\beta}}(\lambda)=\mathbf{0}$ for $\lambda\geq\lambda_{0}$ (plug $\hat{\bm{\beta}}(\lambda_{0})=\mathbf{0}$ in Eq. (7)). The corresponding active and inactive sets are $\mathcal{A}=\operatorname{argmax}_{j\in[d]}|\mathbf{X}^{\top}_{j}\mathbf{y}|$ and $\mathcal{I}=[d]\setminus\mathcal{A}$, respectively, while the equicorrelation sign is $\bm{\eta}_{\mathcal{A}}=\operatorname{sign}(\mathbf{X}^{\top}_{\mathcal{A}}\mathbf{y})$. Since every iteration after the initial setup is identical, assume by induction that the LARS algorithm has computed the first $t+1$ kinks $(\lambda_{0},\hat{\bm{\beta}}(\lambda_{0})),\dots,(\lambda_{t},\hat{\bm{\beta}}(\lambda_{t}))$. At the beginning of the $t$-th iteration, the regularisation parameter is $\lambda=\lambda_{t}$, the Lasso solution is $\hat{\bm{\beta}}(\lambda_{t})$, and the active and inactive sets are $\mathcal{A}$ and $\mathcal{I}$. From the previous solution $\hat{\bm{\beta}}(\lambda_{t})$ we obtain the equicorrelation signs
@@
According to Eq. (8), for $\lambda\leq\lambda_{t}$ sufficiently close to $\lambda_{t}$, the Lasso solution is [^3] From now on, we shall assume $\mathbf{b}=\mathbf{0}$ in order to guarantee the continuity of the Lasso path.
@@
where $\bm{\mu}=\mathbf{X}_{\mathcal{A}}^{+}\mathbf{y}$ and $\bm{\theta}=(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{+}\bm{\eta}_{\mathcal{A}}$. We now decrease $\lambda$ while keeping Eq. (11). As $\lambda$ decreases, two important checks must be made. First, we check when (i.e., we compute the next value of $\lambda$ at which) a variable $i_{t+1}^{\rm join}$ in $\mathcal{I}$ should join the active set $\mathcal{A}$ because it has attain the maximum correlation $|\mathbf{X}_{i_{t+1}^{\rm join}}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})|=\lambda$. We call this event the next joining time $\lambda_{t+1}^{\rm join}$. Second, we check when a variable $i_{t+1}^{\rm cross}$ in the active set $\mathcal{A}$ should leave $\mathcal{A}$ and join $\mathcal{I}$ because $\hat{\beta}_{i_{t+1}^{\rm cross}}$ crossed through zero. We call this event the next crossing time $\lambda_{t+1}^{\rm cross}$.
@@
$\lambda_{t+1}^{\rm join}=\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\},\quad i_{t+1}^{\rm join}=\operatorname*{argmax}_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}.$
@@
$\lambda_{t+1}^{\rm cross}=\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}(\mathcal{A})\},\quad i_{t+1}^{\rm cross}=\operatorname*{argmax}_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}(\mathcal{A})\}.$
@@
1 Input: Vector $\mathbf{y}\in\mathbb{R}^{n}$ and matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ Initialise $\mathcal{A}=\operatorname{argmax}_{j\in[d]}|\mathbf{X}_{j}^{\top}\mathbf{y}|$, $\mathcal{I}=[d]\setminus\mathcal{A}$, $\lambda_{0}=\|\mathbf{X}^{\top}\mathbf{y}\|_{\infty}$, $\hat{\bm{\beta}}(\lambda_{0})=\mathbf{0}$, $t=0$ while $\mathcal{I}\neq\varnothing$ do
       $\bm{\eta}_{\mathcal{A}}\leftarrow\operatorname{sign}\big{(}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\hat{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\big{)}$ $\bm{\mu}\leftarrow\mathbf{X}_{\mathcal{A}}^{+}\mathbf{y}$   // $\bm{\mu}\in\mathbb{R}^{|\mathcal{A}|}$
       $\bm{\theta}\leftarrow(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{+}\bm{\eta}_{\mathcal{A}}$   // $\bm{\theta}\in\mathbb{R}^{|\mathcal{A}|}$
2       $\Lambda_{i}^{\rm cross}\leftarrow\frac{\mu_{i}}{\theta_{i}}\cdot\mathbf{1}\left[\frac{\mu_{i}}{\theta_{i}}\leq\lambda_{t}\right]\quad\forall i\in\mathcal{A}$ $\Lambda_{i}^{\rm join}\leftarrow\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}\quad\forall i\in\mathcal{I}$ $i_{t+1}^{\rm cross}\leftarrow\operatorname{argmax}_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$ and $\lambda_{t+1}^{\rm cross}\leftarrow\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$ $i_{t+1}^{\rm join}\leftarrow\operatorname{argmax}_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}$ and $\lambda_{t+1}^{\rm join}\leftarrow\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}$ $\lambda_{t+1}\leftarrow\max\{\lambda_{t+1}^{\rm join},\lambda_{t+1}^{\rm cross}\}$ $\hat{\bm{\beta}}_{\mathcal{A}}(\lambda_{t+1})\leftarrow\bm{\mu}-\lambda_{t+1}\bm{\theta}$ if $\lambda_{t+1}=\lambda_{t+1}^{\rm cross}$ then
@@ 6      $t\leftarrow t+1$
Output: Coefficients $[(\lambda_{0},\hat{\bm{\beta}}(\lambda_{0})),(\lambda_{1},\hat{\bm{\beta}}(\lambda_{1})),\dots]$
@@
Let us start with the computation of $\mathbf{X}_{\mathcal{A}}^{+}$. Simply computing $\mathbf{X}_{\mathcal{A}}^{+}$ would require $O(n|\mathcal{A}|^{2}+|\mathcal{A}|^{3})$ time. However, the time complexity can be reduced to $O(n|\mathcal{A}|+|\mathcal{A}|^{2})$ by using the Sherman–Morrison formula applied to the Moore-Penrose inverse. More specifically, suppose we have previously computed $\mathbf{X}_{\mathcal{A}}^{+}$ and the index $i_{t+1}^{\rm join}$ is then added to the active set $\mathcal{A}$ to obtain the new active set $\mathcal{A}^{\prime}=\mathcal{A}\cup\{i_{t+1}^{\rm join}\}$. Without lost of generality, write $\mathcal{A}=\{j_{1},\dots,j_{k}\}$ and $\mathcal{A}^{\prime}=\{j_{1},\dots,j_{k},i_{t+1}^{\rm join}\}$, where $k=|\mathcal{A}|$. Let $\mathbf{A}\in\mathbb{R}^{n\times(|\mathcal{A}|+1)}$ be the matrix $\mathbf{X}_{\mathcal{A}}$ augmented with a zero column at the position corresponding to $i_{t+1}^{\rm join}$, i.e., $\mathbf{A}_{l}=\mathbf{X}_{j_{l}}$ if $l\in[k]$, and $\mathbf{A}_{l}=\mathbf{0}$ if $l=k+1$. The augmented matrix is written using a rank-$1$ update as
@@
where $\mathbf{e}_{l}\in\{0,1\}^{|\mathcal{A}|+1}$ is the column vector with $1$ in position $l$ and 00 elsewhere. Therefore, by using the Sherman–Morrison formula applied to the Moore-Penrose inverse (see e.g. [MJ73]), the inverse $\mathbf{X}_{\mathcal{A}^{\prime}}^{+}$ can be computed from the matrix $\mathbf{X}_{\mathcal{A}}^{+}$ and the vectors $\mathbf{X}_{i^{\rm join}_{t+1}}$ and $\mathbf{e}_{k+1}$ using simple matrix and vector multiplication, which requires $O(n|\mathcal{A}|+|\mathcal{A}|^{2})$ time. Regarding the other quantities,
@@
Regarding the number of iterations taken by the LARS algorithm, heuristically, it is on average $O(n)$ [RZ07]. This can be understood as follows. If $n>d$, it would take $O(d)$ steps to add all the variables. If $n<d$, then we only add at most $n$ variables in the case of a unique solution, since $|\mathcal{A}|\leq\min\{n,d\}$. Dropping variables is rare as $\lambda$ is successfully decreased, usually $O(1)$ times. In the worst case, though, the number of iterations can be at most $(3^{d}+1)/2$ as previously mentioned.

### v2

@@ 3.1 Lasso solution
It is known [104, Lemma 14] that there always is a Lasso solution such that $|\mathcal{A}|\leq\min\{n,d\}$. If the Lasso solution is unique, then the active set has size at most $\min\{n,d\}$. The sufficient condition $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$ for uniqueness has appeared several times in the literature [79, 39, 113, 18, 104]. Since the active set $\mathcal{A}$ depends on the Lasso solution at $\mathbf{y},\mathbf{X},\lambda$, the condition that $\mathbf{X}_{\mathcal{A}}$ has full rank is somewhat circular. Because of this, more natural conditions are assumed which imply $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$, e.g., the general position property [88, 28, 30, 104] and the matrix $\mathbf{X}$ being drawn from a continuous probability distribution [104]. We say that a matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ has columns in general position if no $k$-dimensional subspace in $\mathbb{R}^{n}$, for $k<\min\{n,d\}$, contains more than $k+1$ elements from the set {±$\{\pm\mathbf{X}_{1},\dots,\pm\mathbf{X}_{d}\}$, excluding antipodal pairs. It is known [88, 28, 30, 104] that this is enough to guarantee the uniqueness of the Lasso solution.
Fact 21 ([104, Lemma 3]).
If the columns of $\mathbf{X}\in\mathbb{R}^{n\times d}$ are in general position, then the Lasso solution is unique for any $\mathbf{y}\in\mathbb{R}^{n}$ and $\lambda>0$ and is given by Equation 8.
Another sufficient condition was given by Tibshirani [104], who proved the following.
Fact 22 ([104, Lemma 4]).
If the entries of $\mathbf{X}\in\mathbb{R}^{n\times d}$ are drawn from a continuous probability distribution on $\mathbb{R}^{nd}$, then the Lasso solution is unique for any $\mathbf{y}\in\mathbb{R}^{n}$ and $\lambda>0$ and is given by Equation 8 with probability $1$.
It is well known that, when $\operatorname{null}(\mathbf{X}_{\mathcal{A}})=\{\mathbf{0}\}$, the KKT conditions from Fact 19 imply that the regularisation path $\mathcal{P}\triangleq\{\hat{\bm{\beta}}(\lambda)\in\mathbb{R}^{d}:\lambda>0\}$ comprising all the solutions for $\lambda>0$ is continuous piecewise linear, which was first proven by Efron, Hastie, Johnstone, and Tibshirani [32], cf. Equation 2. This fact was later extended to the case when $\operatorname{rank}(\mathbf{X}_{\mathcal{A}})<|\mathcal{A}|$ by Tibshirani [104], who proved that the path of solutions with $\mathbf{b}=\mathbf{0}$ in Equation 6 is continuous and piecewise linear.
Fact 23 ([104, Appendix A]).
@@
$\hat{\bm{\beta}}_{\mathcal{I}}(\lambda)=\mathbf{0}\quad\text{and}\quad\hat{\bm{\beta}}_{\mathcal{A}}(\lambda)=\mathbf{X}_{\mathcal{A}}^{+}(\mathbf{y}-\lambda(\mathbf{X}_{\mathcal{A}}^{+})^{\top}\bm{\eta}_{\mathcal{A}}),$
@@
We call kink a point where the direction $\partial\hat{\bm{\beta}}(\lambda)/\partial\lambda$ changes. The path $\{\hat{\bm{\beta}}(\lambda):\lambda_{t+1}<\lambda\leq\lambda_{t}\}$ between two consecutive kinks $\lambda_{t+1}$ and $\lambda_{t}$ thus defines a linear segment. An important consequence of the above result is that any continuous piecewise linear regularisation path has at most $3^{d}$ linear segments, since two different linear segments must have different equicorrelation signs $\bm{\eta}\in\{-1,0,1\}^{d}$. This result was improved slightly to $(3^{d}+1)/2$ iterations by Mairal and Yu [78], who also exhibited a worst-case Lasso problem with exactly $(3^{d}+1)/2$ number of linear segments.
@@ 3.2 The LARS algorithm
The Lasso solution path $\{\hat{\bm{\beta}}(\lambda):\lambda>0\}$ can be computed by the Least Angle Regression (LARS) algorithm [^2] By LARS algorithm we mean the version of the algorithm that computes the Lasso path and not the version that performs a kind of forward variable selection. which was proposed and named by Efron et al. [32], although similar ideas have already appeared in the work of Osborne et al. [79, 80]. Since its introduction, the LARS algorithm have been improved and generalised [91, 78, 104, 92]. In compressed sensing literature, this algorithm is better known as the Homotopy method [37].
Starting at $\lambda=\infty$, where the Lasso solution is trivially $\mathbf{0}\in\mathbb{R}^{d}$, the LARS algorithm iteratively computes all the kinks in the Lasso solution path by decreasing the parameter $\lambda$ and checking for points where the path’s linear trajectory changes. This is done by making sure that the KKT conditions from Fact 19 are always satisfied. When a kink is reached, a coordinate from the solution $\hat{\bm{\beta}}(\lambda)$ will go from non-zero (being in the active set $\mathcal{A}$) to zero (join the inactive set $\mathcal{I}$), or vice-versa, i.e., from zero (leave $\mathcal{I}$) to non-zero (into $\mathcal{A}$). We shall now describe the LARS algorithm in more details, following [104]. Algorithm 1 summarises the LARS algorithm.
The LARS algorithm starts, without loss of generality, at the first kink $\lambda_{0}=\|\mathbf{X}^{\top}\mathbf{y}\|_{\infty}$, since $\hat{\bm{\beta}}(\lambda)=\mathbf{0}$ for $\lambda\geq\lambda_{0}$ (plug $\hat{\bm{\beta}}(\lambda_{0})=\mathbf{0}$ in Equation 5). The corresponding active and inactive sets are $\mathcal{A}=\operatorname{argmax}_{j\in[d]}|\mathbf{X}^{\top}_{j}\mathbf{y}|$ and $\mathcal{I}=[d]\setminus\mathcal{A}$, respectively, while the equicorrelation sign is $\bm{\eta}_{\mathcal{A}}=\operatorname{sign}(\mathbf{X}^{\top}_{\mathcal{A}}\mathbf{y})$. Since every iteration after the initial setup is identical, assume by induction that the LARS algorithm has computed the first $t+1$ kinks $(\lambda_{0},\hat{\bm{\beta}}(\lambda_{0})),\dots,(\lambda_{t},\hat{\bm{\beta}}(\lambda_{t}))$. At the beginning of the $t$-th iteration, the regularisation parameter is $\lambda=\lambda_{t}$, the Lasso solution is $\hat{\bm{\beta}}(\lambda_{t})$, and the active and inactive sets are $\mathcal{A}$ and $\mathcal{I}$. From the previous solution $\hat{\bm{\beta}}(\lambda_{t})$ we obtain the equicorrelation signs
@@
According to Equation 6, for $\lambda\leq\lambda_{t}$ sufficiently close to $\lambda_{t}$, the Lasso solution is [^3] From now on, we shall assume $\mathbf{b}=\mathbf{0}$ in order to guarantee the continuity of the Lasso path.
@@
where $\bm{\mu}=\mathbf{X}_{\mathcal{A}}^{+}\mathbf{y}$ and $\bm{\theta}=(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{+}\bm{\eta}_{\mathcal{A}}$. We now decrease $\lambda$ while keeping Equation 9. As $\lambda$ decreases, two important checks must be made. First, we check when (i.e., we compute the next value of $\lambda$ at which) a variable $i_{t+1}^{\rm join}$ in $\mathcal{I}$ should join the active set $\mathcal{A}$ because it has attain the maximum correlation $|\mathbf{X}_{i_{t+1}^{\rm join}}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})|=\lambda$. We call this event the next joining time $\lambda_{t+1}^{\rm join}$. Second, we check when a variable $i_{t+1}^{\rm cross}$ in the active set $\mathcal{A}$ should leave $\mathcal{A}$ and join $\mathcal{I}$ because $\hat{\beta}_{i_{t+1}^{\rm cross}}$ crossed through zero. We call this event the next crossing time $\lambda_{t+1}^{\rm cross}$.
@@
$\lambda_{t+1}^{\rm join}=\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\},\qquad i_{t+1}^{\rm join}=\operatorname*{argmax}_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}.$
@@
$\lambda_{t+1}^{\rm cross}=\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}(\mathcal{A})\},\qquad i_{t+1}^{\rm cross}=\operatorname*{argmax}_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}(\mathcal{A})\}.$
@@
1
Input: Vector $\mathbf{y}\in\mathbb{R}^{n}$ and matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$
2
Output: Exact regularisation path $\mathcal{P}$
3
4Initialise $\mathcal{A}=\operatorname{argmax}_{j\in[d]}|\mathbf{X}_{j}^{\top}\mathbf{y}|$, $\mathcal{I}=[d]\setminus\mathcal{A}$, $\lambda_{0}=\|\mathbf{X}^{\top}\mathbf{y}\|_{\infty}$, $\hat{\bm{\beta}}(\lambda_{0})=\mathbf{0}$, $t=0$
5while $\mathcal{I}\neq\varnothing$ do
6       $\bm{\eta}_{\mathcal{A}}\leftarrow\operatorname{sign}\big{(}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\hat{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\big{)}$
      $\bm{\mu}\leftarrow\mathbf{X}_{\mathcal{A}}^{+}\mathbf{y}$
        // $\bm{\mu}\in\mathbb{R}^{|\mathcal{A}|}$
7
      $\bm{\theta}\leftarrow(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{+}\bm{\eta}_{\mathcal{A}}$
        // $\bm{\theta}\in\mathbb{R}^{|\mathcal{A}|}$
8
9      $\Lambda_{i}^{\rm cross}\leftarrow\frac{\mu_{i}}{\theta_{i}}\cdot\mathbf{1}\left[\frac{\mu_{i}}{\theta_{i}}\leq\lambda_{t}\right]\quad\forall i\in\mathcal{A}$
10      $\Lambda_{i}^{\rm join}\leftarrow\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}\quad\forall i\in\mathcal{I}$
11      $i_{t+1}^{\rm cross}\leftarrow\operatorname{argmax}_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$ and $\lambda_{t+1}^{\rm cross}\leftarrow\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$
12      $i_{t+1}^{\rm join}\leftarrow\operatorname{argmax}_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}$ and $\lambda_{t+1}^{\rm join}\leftarrow\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}$
13      $\lambda_{t+1}\leftarrow\max\{\lambda_{t+1}^{\rm join},\lambda_{t+1}^{\rm cross}\}$
14
15      $\hat{\bm{\beta}}_{\mathcal{A}}(\lambda_{t+1})\leftarrow\bm{\mu}-\lambda_{t+1}\bm{\theta}$
16
17      if $\lambda_{t+1}=\lambda_{t+1}^{\rm cross}$ then
18
@@ 24      $t\leftarrow t+1$
return coefficients $[(\lambda_{0},\hat{\bm{\beta}}(\lambda_{0})),(\lambda_{1},\hat{\bm{\beta}}(\lambda_{1})),\dots]$
@@
Let us start with the computation of $\mathbf{X}_{\mathcal{A}}^{+}$. Simply computing $\mathbf{X}_{\mathcal{A}}^{+}$ would require $O(n|\mathcal{A}|^{2}+|\mathcal{A}|^{3})$ time. However, the time complexity can be reduced to $O(n|\mathcal{A}|+|\mathcal{A}|^{2})$ by using the Sherman–Morrison formula applied to the Moore-Penrose inverse. More specifically, suppose we have previously computed $\mathbf{X}_{\mathcal{A}}^{+}$ and the index $i_{t+1}^{\rm join}$ is then added to the active set $\mathcal{A}$ to obtain the new active set $\mathcal{A}^{\prime}=\mathcal{A}\cup\{i_{t+1}^{\rm join}\}$. Without lost of generality, write $\mathcal{A}=\{j_{1},\dots,j_{k}\}$ and $\mathcal{A}^{\prime}=\{j_{1},\dots,j_{k},i_{t+1}^{\rm join}\}$, where $k=|\mathcal{A}|$. Let $\mathbf{A}\in\mathbb{R}^{n\times(|\mathcal{A}|+1)}$ be the matrix $\mathbf{X}_{\mathcal{A}}$ augmented with a zero column at the position corresponding to $i_{t+1}^{\rm join}$, i.e., $\mathbf{A}_{\ell}=\mathbf{X}_{j_{\ell}}$ if $\ell\in[k]$, and $\mathbf{A}_{\ell}=\mathbf{0}$ if $\ell=k+1$. The augmented matrix is written using a rank-$1$ update as
@@
where $\mathbf{e}_{\ell}\in\{0,1\}^{|\mathcal{A}|+1}$ is the column vector with $1$ in position $\ell$ and 00 elsewhere. Therefore, by using the Sherman–Morrison formula applied to the Moore-Penrose inverse (see [73]), the inverse $\mathbf{X}_{\mathcal{A}^{\prime}}^{+}$ can be computed from the matrix $\mathbf{X}_{\mathcal{A}}^{+}$ and the vectors $\mathbf{X}_{i^{\rm join}_{t+1}}$ and $\mathbf{e}_{k+1}$ using simple matrix and vector multiplication, which requires $O(n|\mathcal{A}|+|\mathcal{A}|^{2})$ time. Regarding the other quantities,
@@
Regarding the number of iterations taken by the LARS algorithm, heuristically, it is on average $O(n)$ [91]. This can be understood as follows. If $n>d$, it would take $O(d)$ steps to add all the variables. If $n<d$, then we only add at most $n$ variables in the case of a unique solution, since $|\mathcal{A}|\leq\min\{n,d\}$. Dropping variables is rare as $\lambda$ is successfully decreased, usually $O(1)$ times. In the worst case, though, the number of iterations can be at most $(3^{d}+1)/2$ as previously mentioned.